Operations Report – May 3rd

Crew 235 Operations Report 3-MAY-2021

SOL: 1

Name of person filing report: Atila Meszaros

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 0

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 32% at 7:01 pm (Before generator is run at night)

Notes on power system: Yesterday, the SOC went nominal at 13:00 after a couple of days in 0%. It jumped to 100%, and it has been slowly decreasing since then. This morning it got to a max of 58% and it is decreasing again. We have had power through the station all this time and there haven’t been any variations on the power, but the system is definitely showing the opposite.

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 74% 2. 80%

Water (loft tank): 25 gallons

Water Meter: 150932.6 units

Water (static tank): 25 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 207.8

Beginning charge: 100%

Ending charge: 96%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: Training and EVA #1

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 0

Notes on ATVs: ATV’s will be filled with Free Ethanol Gas to be ready for tomorrow’s EVA.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging. We have 4 fully operational suits, but I’d love to get a couple more working.

Summary of Hab operations: Spaceward Bound Crew went into operations and sim training.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Start of the microbiology experiment.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Thank you for being with us tonight!